Use the interactive Punnet Squares Level 2 to answer the following questions. Suppose the parents indicated in the interactive produced 208 peas. Determine the expected number of yellow and wrinkled offspring Number Now suppose one of the pea plants is homozygous recessive for both genes, resulting in the cros Yy Rrxyy IT. Calculate the percentage of yellow and wrinkled offspring Number Suppose one of the pea plants is homozygous recessive for seed color, resulting in the cross Yy Rr xyy Rr. Which of the following represents the ratio of phenotypes of the offspring? O 9 yellow, round 3 yellow, wrinkled: 3 green, round 1 green, wrinkled O 4 yellow, round: 4 yellow, wrinkled :4 green, round: 4 green, wrinkled O 6 yellow, round 6 yellow, wrinkled 2 green, round: 2 green, wrinkled O 6 yellow, round 2 yellow, wrinkled 6 green, round 2 green, wrinkled
Answer #1

1.

Answer :- 39

Explanation :-

Dihybrid Ratio of cross between Yellow color,green color, Round seed and wrinkled seed - 9:3:3:1

cross between - YyRr × YyRr produces :-

Yellow Round seed = 9/16

Green Round seed = 3/16

Yellow wrinkled seed = 3/16

Green wrinkled seed = 1/16

In this way , if total pea is 208 ,number of yellow and wrinkled offspring will be 39.

2.

Answer is :- 25 %

Explanation :-

If one of pea plant is homozygous recessive for both genes and it results in cross Yy Rr × yy rr, then the percentage of yellow and wrinkled offspring will be 25%.

3.

Answer is :- 6 yellow, round : 2 yellow, wrinkled : 6 green, round : 2 green, wrinkled

Explanation :-

one of the pea plants is homozygous recessive for seed color, resulting in the cross Yy Rr × yy Rr.

6 yellow, round : 2 yellow, wrinkled : 6 green, round : 2 green, wrinkled , represents the ratio of phenotypes of the offspring.

(A) is wrong answer because it has written 9 yellow, round: 3 yellow, wrinkled: 3 green, round: 1 green, wrinkled.

(B) is wrong answer because it has written 4 yellow, round: 4 yellow, wrinkled : 4 green, round : 4 green, wrinkled.

(C) is wrong answer because it has written 6 yellow, round: 6 yellow, wrinkled: 2 green, round: 2 green, wrinkled.


answered by: ANURANJAN SARSAM
Answer #2

1.

Answer :- 39

Explanation :-

Dihybrid Ratio of cross between Yellow color,green color, Round seed and wrinkled seed - 9:3:3:1

cross between - YyRr × YyRr produces :-

Yellow Round seed = 9/16

Green Round seed = 3/16

Yellow wrinkled seed = 3/16

Green wrinkled seed = 1/16

In this way , if total pea is 208 ,number of yellow and wrinkled offspring will be 39.

2.

Answer is :- 25 %

Explanation :-

If one of pea plant is homozygous recessive for both genes and it results in cross Yy Rr × yy rr, then the percentage of yellow and wrinkled offspring will be 25%.

3.

Answer is :- 6 yellow, round : 2 yellow, wrinkled : 6 green, round : 2 green, wrinkled

Explanation :-

one of the pea plants is homozygous recessive for seed color, resulting in the cross Yy Rr × yy Rr.

6 yellow, round : 2 yellow, wrinkled : 6 green, round : 2 green, wrinkled , represents the ratio of phenotypes of the offspring.

(A) is wrong answer because it has written 9 yellow, round: 3 yellow, wrinkled: 3 green, round: 1 green, wrinkled.

(B) is wrong answer because it has written 4 yellow, round: 4 yellow, wrinkled : 4 green, round : 4 green, wrinkled.

(C) is wrong answer because it has written 6 yellow, round: 6 yellow, wrinkled: 2 green, round: 2 green, wrinkled.

*If you have any doubt write in comment box and i will solve it and if you are satisfied plz rate my answer...